OpenStudy (anonymous):

PLEASE EMERGENCY PLEASE!!!!!!!!!!! Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown. line through the points negative 3 comma negative 4 and negative 1 comma 3 with the point negative 2 comma 2 not on the line y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (anonymous):

points are (-3,-4) & (-1,3) equation of line through (-3,-4) & (-1,3) is \[y-(-4)=\frac{ 3+4 }{-1+3 }(x-(-3))\] \[y+4=\frac{ 7 }{ 2 }(x+3)\] 2y+8=7x+21 2y=7x+21-8 2y=7x+13 ....(1) Any line parallel to (1) is 2y=7x+c .....(2) If it passes through (-2,2) 2*2=7*-2+c c=4+14=18 (2) is the required equation ,where c is any real numbers except 18